OpenStudy (anonymous):

10x+6=2(6x-4) x=7ok i show my work and that expand out terms on the right hand side which is 10x+6=12x-8 add (-12x-6) to both sides -2x=-14 divide both sids by -2 which gave me x=7 but i need to check my work please help with checking my work

jhonyy9 (jhonyy9):

10*7 +6=2(6*7-4) 70 +6=84-8 76=76

OpenStudy (anonymous):

thank you

jhonyy9 (jhonyy9):

yw good luck bye